The following stem-leaf plot represents the age data of a sample of size 40 selected from all faculty members of a local high school, where “4 3” represents 43 years old.

2 122267
3 1234567899
4 123456789
5 123455
6 013556789

Calculate the five-number summary, and find IQR.
1
Expert's answer
2020-12-22T20:13:22-0500

21 22 22 22 26 27 31 22 33 34

35 36 37 38 39 39 41 42 43 44

45 46 47 48 49 51 52 53 54 55

55 60 61 63 65 65 66 67 68 69

The five-number summary:

Minimum =21

Maximum = 69

First quartile

"Q_1 = \frac{1}{4}(n+1)^{th} \\

= \frac{41}{4} \\

= 10.25^{th}

10^{th} + 0.25(11^{th}-10^{th}) \\

= 34+0.25(35-34) \\

Q_1 = 34.25 \\

Q_2 = \frac{2}{4}(n+1)^{th} \\

= \frac{41^{th}}{2} \\

= 20.5^{th} \\

= 20^{th} + 0.5(21^{th}-20^{th}) \\

= 44 + 0.5(45-44) \\

Q_2 = 44.05 \\

Median = Q_2 = 44.05 \\

Q_3 = \frac{3}{4}(n+1)^{th} \\

= 30.75^{th} \\

= 55 + 0.75(55-55) \\

= 55 \\

Q_3 = 55 \\

IQR = Q_3-Q_1 \\

= 55 -34.25 \\

= 20.75"